HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ACCURATE FIRST DISSOCIATION LIMIT OF THE OZONE MOLECULE FROM AB INITIO CALCULATIONS

OLEG EGOROV, ROMAN V. KOCHANOV, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Tomsk State University 36, Lenin Ave., Tomsk,

634050 Russia; \underline{VLADIMIR TYUTEREV} , Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia; VIATCHESLAV KOKOOULINE, University of Central Florida, Orlando, USA

In this work a good agreement was achieved between the theoretical and the empirical value of the first dissociation limit (D_{e}) of O_{3}(X̃^{1}A_{1}). The ab initio calculations were done with the extended CAS(24, 15) active space and with the inclusion of the spin-orbit coupling between the nine electronic states approaching in D_{e}. Finally, the full-dimensional potential energy surface was developed in the scattering (Jacobi) coordinates for the long-range region (R_{O_{2}-O} ≥ 5 a_{0}).

The study was supported by the RSF grant no. 19-12-00171-P.

p-number: p113  Submitted on Wed, 14 Jun 2023 08:03:16 +0200